OpenStudy (anonymous):

The lengh of a rectangular garden is 7ft longer than itgs width. The perimeter of the garden is 38ft. Find the width and length of the garden. please help last question on my side guide

OpenStudy (anonymous):

If w is the garden's width, then the length is w+7 Solve the following equation for w: 38 = 2 (w + (w + 7)) w = 6 w+7 = 13